Hoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singapore. The company has determined that if a truck is driven 174,000 kilometers during a year, the average operating cost is 13.8 cents per kilometer. If a truck is driven only 116,000 kilometers during a year, the average operating cost increases to 17.7 cents per kilometer. Required: 1. Using the high-low method, estimate the variable and fixed cost elements of the annual cost of the truck operation. (Do not round your intermediate calculations. Round the Variable cost per kilometer to 3 decimal places and Fixed cost answer to nearest whole dollar amount.) 2. Express the variable and fixed costs in the form Y = a + bX. (Do not round your intermediate calculations. Round the Variable cost per kilometer to 3 decimal places.) 3. If a truck were driven 145,000 kilometers during a year, what total cost would you expect to be incurred? (Do not round intermediate calculations.)
Answer #1

1.

Truck is driven 174,000 kilometers during a year, the average operating cost is 13.8 cents per kilometer.

so operating cost = 174,000 x 13.8/100

= $24,012

Truck is driven only 116,000 kilometers during a year, the average operating cost increases to 17.7 cents per kilometer.

So Operating cost = 116,000 x 17.7/100

= $20,532

Variable cost per unit = (Highest activity cost - Lowest activity cost)/(Highest activity - Lowest activity)

= (24,012-20,532)/(174,000-116,000)

= 3,480/58,000

= 0.06 per kilometer

Fixed cost = Highest activity cost - Highest activity x Variable cost per kilometer

= 24,012 - 174,000 x 0.06

= $13,572

2.

Y = a + bX

= 0.06 + 13,572X

3. Truck were driven 145,000 kilometers during a year, Total cost would you expect to be incurred = Fixed cost + Variable cost per kilometer x Activity Kilometer

= 13,572 + 0.06 x 145,000

= $22,272

Kindly comment if you need further assistance. Thanks‼!